Sstava:

Carlos has box of sticks. 2/5 are raspberry of the remaining 1/3 are grape and rest are cherry. If there are 48 raspberry and cherry altogether how many total sticks

jhonyy9:

so i can help you than you will be online - please tag me here

Mercury:

let's let total # of sticks = x raspberry = (2/5)x "of the remaining, 1/3 are grape), so (1 - 2/5)x remains, and grape = (1/3)(1 - 2/5)x "the rest are cherry" so cherry = x - (2/5)x - [(1/3)(1 - 2/5)x] "48 raspberry and cherry altogether" ---> (2/5)x + x - (2/5)x - [(1/3)(1 - 2/5)x] = 48, solve for x